# BoxingGym: Benchmarking Progress in Automated Experimental Design and Model Discovery

### v1

@@ Appendix A Full Results
See Fig. 4 for the performance of the two baseline agents on peregrines, hyperbolic discounting, and IRT environments for different number of experiments. See Tab. 3 for the EI regrets of the two agents. See Tab. 4, and Tab. 5 for prediction errors across all environments.
Figure 4: EIG regrets and standardized errors compared. We plot the standardized errors (top row) and the EIG regrets (bottom row) for the two agents, gpt-4o (solid line) and Box’s Apprentice (dashed line) across three domains: Peregrines (left), Hyperbolic Discounting (center) and IRT (right). Error bars are 95% CIs across 5 runs.
Env Goal EI Regret (gpt-4o) EI Regret (box’s apprentice)
Hyperbolic Discounting Choice 0.57 / 0.61 0.55 / 0.62
Discount 0.69 / - - / -
Location Finding Signal 15.3 / 11.8 12.6 / 15.3
Source Location 16.8 / - - / -
Death Process Num Infected 0.037 / 0.042 0.029 / 0.019
Infection Rate 0.108 / - - / -
IRT Correctness 0.035 / 0.031 0.031 / 0.033
Dugongs Length 0.20 / 0.17 0.19 / 0.20
Peregrines Population 0.26 / 0.38 0.25 / 0.66
Mastectomy Survival 0.084 / 0.082 0.079 / 0.075
Predator-Prey Population - / - - / -
Emotions Prediction 0.538 / - - / -
Moral Machines Judgement 0.046 / - 0.045 / -
Table 3: EI Regrets for GPT-4o and Box’s Apprentice Across Different Tasks. EI regrets for prior and no prior conditinos are separated by ‘/’.
@@
Hyperbolic Discounting Choice 0.31$\pm$0.18 0.96$\pm$0.14 0.74$\pm$0.21 0.95$\pm$0.07 0.74$\pm$0.14 1.0$\pm$0.00
Hyperbolic Discounting Discount -$0$$0$
Location Finding Signal 0.96$\pm$0.58 1.17$\pm$0.60 1.24$\pm$0.96 0.5$\pm$0.54 0.97$\pm$0.72 0.63$\pm$0.71
Location Finding Source Location 1.29$\pm$1.3 - -0.15 $\pm$0.4 - - -
Death Process Num Infected 1.19$\pm$1.09 0.19$\pm$0.96 0.46$\pm$0.76 0.74$\pm$1.14 0.75$\pm$0.75 1.61$\pm$1.60
Death Process Infection Rate 0.13$\pm$0.37 - 1.64$\pm$1.12 - - -
IRT Correctness $0$0.16$\pm$0.26 0$\pm$0.11 0.08$\pm$0.32 -0.28$\pm$0.26 -0.16$\pm$0.20
Dugongs Length 0.06$\pm$0.12 -0.02$\pm$0.04 -$0$$0$0.08$\pm$0.01 -0.08$\pm$0.01
Peregrines Population 2.29$\pm$1.20 2.21$\pm$1.57 -0.65$\pm$0.03 -0.67$\pm$0.01 -0.63$\pm$0.06 -0.66$\pm$0.02
Mastectomy Survival 0.18$\pm$0.37 0.00$\pm$0.28 0.27$\pm$0.19 0.36$\pm$0.27 1.00$\pm$0.27 0.21$\pm$0.16
Predator-Prey Population 0.08$\pm$0.09 0.73$\pm$0.05 -0.45$\pm$0.02 -0.43$\pm$0.02 -0.26$\pm$0.16 -0.40$\pm$0.03
Emotions Prediction 0.74$\pm$0.29 - 0.82$\pm$0.34 - 0.87$\pm$0.35 -
Moral Machines Judgement 0.32$\pm$0.26 - 0.44$\pm$0.16 - 0.60$\pm$0.13 -
Table 4: Performance of GPT-4o Across Different Tasks. Numbers shown are standardized errors. Errors with prior (top line) and without prior (bottom line) appear on different lines. Errors are averaged across 5 runs.
@@
Hyperbolic Discounting Choice 0.66 $\pm$ 0.25 0.66 $\pm$ 0.25 1.17 $\pm$ 0.14 0.91 $\pm$ 0.09 0.66 $\pm$ 0.30 0.74 $\pm$ 0.42
Location Finding Signal 0.99 $\pm$ 0.58 1.18 $\pm$ 0.64 1.45 $\pm$ 1.60 0.83 $\pm$ 0.600 1.18 $\pm$ 1.12 -0.01 $\pm$ 0.30
Death Process Num Infected 3.79 $\pm$ 1.68 -0.90 $\pm$ 0.05 -1.02 $\pm$ 0.05 -0.61 $\pm$ 0.30 0.58 $\pm$ 0.85 0.50 $\pm$ 1.26
IRT Correctness 0.44 $\pm$ 0.36 $0.12\pm 0.24$ $-0.12\pm 0.14$ $0.12\pm 0.14$0.08 $\pm$ 0.39 0.2 $\pm$ 0.40
Dugongs Length 0.26 $\pm$ 0.12 0.05 $\pm$ 0.10 -0.08 $\pm$ 0.02 -0.09 $\pm$ 0.004 $-0.09\pm 0.005$ $-0.08\pm 0.004$
Peregrines Population 2.71 $\pm$ 0.60 1.62 $\pm$ 0.47 0.04 $\pm$ 0.21 0.95 $\pm$ 0.86 0.97 $\pm$ 1.38 -0.19 $\pm$ 0.79
Mastectomy Survival 0.14 $\pm$ 0.41 0.73 $\pm$ 0.15 0.55 $\pm$ 0.24 0.64 $\pm$ 0.15 0.91 $\pm$ 0.28 0.27 $\pm$ 0.23
Moral Machines Judgement 0.97 $\pm$±i 0.33 0.89 $\pm$ 0.21 0.56 $\pm$ 0.18
@@ D.2 Hyperbolic Discounting
The hyperbolic discounting domain has two hidden variables $(k,\alpha)$ to describe a participant’s behavior, where each participant is asked to choose between an immediate reward \$$\$iR$ or a delayed reward \$$\$dR$ in $D$ days. The experiment is outlined in table 7 below.
@@
where the HalfNormal distribution is a normal distribution truncated at 0. For each test, there are three variables in design: $iR$, $dR$, and $D$. We give values to each choice: receiving the immediate reward \$$\$iR$ has value $V_{i}=iR$, while receiving the delayed reward \$$\$dR$ in $D$ days has value $V_{d}=\frac{dR}{1+kD}$. Then, whether each participant’s chooses the delayed reward in each scenario is characterized as a Bernoulli random variable $X\sim Bernoulli(p)$ where the probability of choosing the delayed reward is given by
@@ D.6 Peregrines
In this environment, the population count of peregrine falcons at time $t$ is modeled using a Poisson regression model with parameters $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ The time values range between 0 and 5. The population count $C$ at a given time $t$ is generated from a Poisson distribution with a mean that is a function of $t$ and the parameters $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$. The function representing the log of the mean population count $\lambda$ is defined as:
@@ D.7 Survival Analysis: Mastectomy
[…]

### v2

@@ Appendix A Full Results
See Tab. 1, and Tab. 2 for  [^3] We omit the predatory-prey and Emotions domains for Box’s Apprentice, since GPT-4o could not reliably produce pymc programs prediction errors across all environments for GPT-4o and the Box’s apprentice with GPT-4o. Full results are available in the Github Repository.
@@
Hyperbolic Discounting Choice
[…]
Hyperbolic Discounting Discount
[…]
Location Finding Signal
[…]
Location Finding Source Location
[…]
-0.15 $\pm$0.4
Death Process Num Infected
[…]
Death Process Infection Rate
[…]
IRT Correctness
[…]
$0±\pm0$.13
[…]
Dugongs Length
[…]
Peregrines Population
[…]
-$6±\pm0$.03
Mastectomy Survival
[…]
Predator-Prey Population
[…]
Emotions Prediction
[…]
N/A
[…]
N/A
[…]
N/A
Moral Machines Judgement
[…]
N/A
[…]
N/A
[…]
N/A
Table 1: Performance of GPT-4o Across Different Tasks. Numbers shown are normalized-0 errors. Errors with prior (top line) and without prior (bottom line) appear on different lines. Errors are averaged across 5 runs.
@@
Hyperbolic Discounting Choice
0.66 $\pm$ 0.25
0.66 $\pm$ 0.25
1.17 $\pm$ 0.14
0.91 $\pm$ 0.09
0.66 $\pm$ 0.30
0.74 $\pm$ 0.42
Location Finding Signal
0.99 $\pm$ 0.58
1.18 $\pm$ 0.64
1.45 $\pm$ 1.60
0.83 $\pm$ 0.60
1.18 $\pm$ 1.12
-0.01 $\pm$ 0.30
Death Process Num Infected
3.79 $\pm$ 1.68
-0.90 $\pm$ 0.05
-1.02 $\pm$ 0.05
-0.61 $\pm$ 0.30
0.58 $\pm$ 0.85
0.50 $\pm$ 1.26
IRT Correctness
0.44 $\pm$ 0.36
$0.12\pm 0.24$
$0.12\pm 0.14$
$0.12\pm 0.14$
-0.08 $\pm$ 0.39
0.2 $\pm$ 0.40
Dugongs Length
0.26 $\pm$ 0.12
0.05 $\pm$ 0.10
-0.08 $\pm$ 0.02
-0.09 $\pm$ 0.004
$0.09\pm 0.005$
$0.08\pm 0.004$
Peregrines Population
2.71 $\pm$ 0.60
1.62 $\pm$ 0.47
0.04 $\pm$ 0.21
0.95 $\pm$ 0.86
0.97 $\pm$ 1.38
-0.19 $\pm$ 0.79
Mastectomy Survival
0.14 $\pm$ 0.41
0.73 $\pm$ 0.15
0.55 $\pm$ 0.24
0.64 $\pm$ 0.15
0.91 $\pm$ 0.28
0.27 $\pm$ 0.23
Moral Machines Judgement
0.97 $\pm$ 0.33
0.89 $\pm$ 0.21
0.56 $\pm$ 0.18
@@ Appendix B LLM Agent
Models were configured with a temperature parameter of 0.0 to ensure deterministic outputs. Maximum token limits were set to 512 tokens for instruct models and 1024 tokens for thinking variants, providing sufficient thinking tokens for generating an answer without multiple retries.
GPT-4o and Claude-3.7-Sonnet were accessed via their APIs, while all other models were deployed using vLLM. For the vLLM-served models, we utilized a dual A40 GPU configuration: one GPU dedicated to model serving and the other for inference execution through the vLLM endpoint. This architecture ensured optimal resource allocation and performance stability throughout the experimental process.
Each OED experimental run consisted of 10 predictions conducted after 0, 1, 3, 5, 7, and 10 observations, respectively. Comprehensive log files were generated for each set of predictions to facilitate subsequent analysis. Execution time varied across model architectures, with most configurations requiring approximately 2-3 minutes per run (defined as a single seed, configuration, and environment combination). Models accessed through external APIs typically required longer execution times due to network latency and rate limiting considerations. Discovery experiments reduced execution times compared to OED experiments due to the decreased number of required API calls.
@@ D.2 Hyperbolic Discounting
The hyperbolic discounting domain has two hidden variables $(k,\alpha)$ to describe a participant’s behavior, where each participant is asked to choose between an immediate reward $\mathdollar iR$ or a delayed reward $\mathdollar dR$ in $D$ days. The experiment is outlined in table 4 below.
@@
where the HalfNormal distribution is a normal distribution truncated at 0. For each test, there are three variables in design: $iR$, dRdR, and $D$. We give values to each choice: receiving the immediate reward $\mathdollar iR$ has value $V_{i}=iR$, while receiving the delayed reward $\mathdollar dR$ in $D$ days has value $V_{d}=\frac{dR}{1+kD}$. Then, whether each participant’s chooses the delayed reward in each scenario is characterized as a Bernoulli random variable $X\sim Bernoulli(p)$ where the probability of choosing the delayed reward is given by
@@ D.6 Peregrines
In this environment, the population count of peregrine falcons at time $t$ is modeled using a Poisson regression model with parameters $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ . The time values range between 0 and 5. The population count $C$ at a given time $t$ is generated from a Poisson distribution with a mean that is a function of $t$ and the parameters $\alpha$, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$. The function representing the log of the mean population count $\lambda$ is defined as:
@@ D.7 Survival Analysis: Mastectomy
$\text{death\_outcome}\sim\text{Bernoulli}(p(\text{death}))$ (14)